OpenStudy (anonymous):

Choose the scientific notation for 800. 1. 8 x 10^3 2. 8 x 10^2 3. 8 x 10^1 4. 8 x 10^0

OpenStudy (anonymous):

800=8*10^2......10^2 =10*10=100.....so 100*8=800